OpenStudy (anonymous):

Complete the table using the given values of x and the equation 3x – y = 8. Show your work. x y 0 1 2 3 4

OpenStudy (anonymous):

0: 3(0)-y=8 0-y=8 y=-8 1: 3(1)-y=8 3-y=8 y=-5 and so on and so forth

OpenStudy (anonymous):

first rearrange the equation as follows: 3x - y = 8 3x - 8 = y

OpenStudy (anonymous):

or, you could put 3x-7=8 into standard form (3x-8=y) and for 0, use the y intercept (-8) and add 3 to the previous y value because it is the slope.

OpenStudy (anonymous):

now substitute the given values of x one by one and get the corresponding values of y... eg first value of x is 0 so 3(0) - 8 = y ==> -8 = y

OpenStudy (anonymous):

so you would have: x y 0 -8 1 -5 2 -2 3 1 4 4 5 7 6 10